OpenStudy (anonymous):

root of 1+cosA/1-cosA PROVE IT

OpenStudy (aravindg):

\[\large \sqrt{\frac{1+\cos A}{1-\cos A}}=?\]

OpenStudy (anonymous):

dey didn't giv dey tld jst 2 prove

OpenStudy (aravindg):

what you mean to prove?

OpenStudy (aravindg):

for proving we need both lhs and rhs

OpenStudy (anonymous):

i thnk it will b cosecA+cotA

OpenStudy (aravindg):

anyway the expressions value is \[\large \cot \theta/2\]

OpenStudy (anonymous):

cn u jst solve it and show

OpenStudy (aravindg):